# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 4 Continual learning strategy
Owing to frequently encountered domain shifts in medical applications, CL strategies aim to tackle performance drop due to domain shifts and thus might become a necessity in clinical applications Pianykh et al. (2020); Lee and Lee (2020). All the available strategies for preventing catastrophic forgetting can be broadly categorized into three main categories: (a) rehearsal-based approaches where a small memory is used to store previous episode data in some form, (b) regularization-based methods where the aim is to control weight update to minimize forgetting the previous learning, (c) architectural-based methods which mainly aim to keep some network parameters isolated for each episode, and (d) hybrid category which offers various combination of any of these three categories. Figure 6 presents a taxonomy of CL techniques and their adaptation in the medical field by various works. Further, the sub-technique category for each work is also mentioned in the figure. Additionally, a pie chart showing the ratio of works falling in the above-mentioned categories is shown via Figure 5. It indicates the popularity of using one or more strategies together in hybrid settings owing to better classification or segmentation performances in sequential learning. Now, we discuss each category in detail in the following sections.
@@
Reference (year) CL scenario & application CL technique
Hofmanninger et al. (2020a) Chest CT classification: DIS (3 Epi.) Experience replay with a dynamic memory size where samples are ranked by computing Gram matrix on last activation maps
Srivastava et al. (2021) Chest x-ray classification: DIS(3 Epi.) Latent replay
Perkonigg et al. (2021) Brain age estimation: DIS (3 Epi.) Experience replay
Perkonigg et al. (2022) DIS: three separate experiments for 3 applications: cardiac segmentation (4 Epi.), lung nodule detection (4 Epi.), and brain age estimation (4 Epi.) Experience replay
Karthik et al. (2022) DIS on MRI data for brain sclerosis lesions segmentation (8 Epi.) Random samples from each episode kept for experience replay
Li et al. (2022) DIS: CMR based Cardiac segmentation across 4 scanners Generative replay: style-oriented replay
Bera et al. (2023) TIS (binary segmentation of prostate, spleen, and hippocampus), DIS (binary hippocampus segmentation across 2 datasets), DIS (binary prostate segmentation across 4 datasets) Experience replay based rehearsal strategy: store max K exemplars (half of the exemplars based on fraction of positive class content inside the image and other half based on gradient variation)
Wang et al. (2023b) (a) Segmentation for endoscopy: CIS (2-3 Epi.) (b) surgical instrument segmentation (2 Epi.) Generative replay
@@ 4.1.2 Latent replay-based
While experience replay requires storing the past samples in raw form, it can cause serious privacy violations in critical medical applications. Therefore, a less concerning direction, i.e., storing features instead of raw images, is also explored. For continual chest x-ray classification application in domain shift conditions, Srivastava et al. (2021) explore leveraging vector-quantization to store and replay hidden representations under memory constraints.
@@ 4.1.4 Memory buffer optimization
Experience replay-based approaches might overfit the stored sub-samples and seem to be bounded by joint training. Alternatively, memory buffer optimization-based approaches following constrained optimization provide solutions that have more scope for both plasticity and stability. The core idea is to perform gradient projection. For example, the Gradient Episodic Memory (GEM) Lopez-Paz and Ranzato (2017) approach corrects the gradient computed on a mini-batch during stochastic gradient descent by utilizing an exemplar set. This prevents changes that could degrade the performance of the network on the exemplar set. Further, the A-GEM approach Chaudhry et al. (2018a) relaxes the projection in one direction computed by randomly selected samples from the memory buffer where both the buffer and sampling size are tunable hyper-parameters. There are some early attempts in this area in the medical realm. For example, for the fundus disease diagnosis problem, Shu et al. (2022) presents a gradient regularisation approach for preventing forgetting and a replay-oriented consistency calculation method combined with a subspace weighting strategy to promote model adaptability. The authors employ gradient regularisation in conjunction with a replay-oriented technique. The replay-oriented technique adaptively improves the update for incremental domains without carefully choosing exemplars for replay, whereas gradient regularisation preserves information from previous ones.
@@
Reference (year) CL scenario & application CL technique
Baweja et al. (2018) MRI based normal brain structures segmentation: TIS (2 Ep.) EWC
McClure et al. (2018) Axial and sagittal brain segmentation DIS (4 Ep.) Distributed Weight Consolidation (DWC)
van Garderen et al. (2019) MR data based glioma segmentation: DIS (2 epi) EWC
Özgün et al. (2020) Brain MRI segmentation: DIS (4 epi) Learning rate regularization built over memory aware synapses technique
Chen and Tang (2022) CIS for histopathology-based breast cancer classification (4 Epi) Regularization: EWC
Zhang et al.; Zhang et al. (2023d; 2021) DIS for (a) MRI based prostate segmentation, (b) DIS for retinal image based optic cup and disc segmentation Shape and semantics-based selective regularization to penalize changes of parameters with high joint shape and semantics-based importance
Zhu et al. (2023) (a) DIS: binary prostate segmentation from T2-weighted MRI scans collected from 6 sites, (b) cross-site and cross-modality 4-class abdominal segmentation between CT and MRI scans, (c) same-site cross-modality muscle segmentation of muscles and bones between MRI and mDixon scans Knowledge distillation
Liu et al. (2023a) Incremental brain tumor segmentation: CIS with domain shift (3 epi) Pseudo label-based knowledge distillation
@@ 4.2.1 Prior-focused regularization
Apart from EWC and SI, various other methods have been contributed. Popularly, the Memory Aware Synapses (MAS) approach Aljundi et al. (2017) calculates the importance of weights with a model of Hebbian learning in the biological system, which relies on the sensitivity of the output function and can hence be utilized in an unsupervised manner. For medical segmentation, a MAS-inspired learning rate regularization approach was contributed by Özgün et al. (2020) for sequential training on different domains. It involves reducing the learning rate for important parameters to prevent forgetting, offering a more direct approach compared to the surrogate loss used in MAS.
@@ 4.2.2 Data-focused regularization
The data-focused regularization method Li and Hoiem (2017) aims to distill knowledge from a model trained on the previous tasks to the model trained on the new task in order to consolidate previously learned knowledge. Typically, the previous model acts as a teacher and the current model as a student while adopting knowledge-distillation Gou et al. (2021) to avoid catastrophic forgetting. Usually, all the old data should be available in knowledge distillation, which is not the case with CL; therefore, as an alternative, few old data or the current data or the generated data are explored for the same. The loss function has an additional distillation loss for replayed data. Each input is replayed with a soft target obtained using the stored model.
Silver and Mercer (2002) first proposed to use previous task model outputs given new task input images, mainly for improving new task performance. Later, in natural image classification application, Li and Hoiem (2017) re-introduced the concept as LwF technique. A copy of the previous model parameters is stored before learning the new task, and then it is used to get the soft labels for the new task as the target for the classifiers of the previous tasks. The available ground truth is used as the target for the new task classifier. LwF has been adopted in various medical applications like incremental brain MRI segmentation Ozdemir et al. (2018), chest x-ray classification Lenga et al. (2020), disease classification Derakhshani et al. (2022), histopathology tumor classification Kaustaban et al. (2022), etc.
Other works in the natural image processing domain Jung et al. (2016); Zhang et al. (2020) have been introduced with LwF-related ideas; however, it has been shown that this strategy is vulnerable to domain shift between tasks Aljundi et al. (2017). It was pointed out that LwF would not result in good performance if the data distributions between different tasks are quite diverse Rannen et al. (2017). To overcome this, Rannen et al. (2017) facilitate incremental integration of shallow auto-encoders to constrain task features in their corresponding learned low dimensional space. Hence, they further trained an autoencoder for each task to learn the most important features corresponding to the task and used it to preserve knowledge.
@@
Reference (year) CL scenario & application CL technique
@@
Bayasi et al. (2021) Skin lesion image classification: DIS (6 Ep.) Architecture based (fixed but partitioned network)
@@
González et al. (2022) DIS with (2 Ep.) for MRI based Hippocampus Segmentation Architecture based approach based on out of distribution detection concept: maintain multivariate Gaussians for all past learning created on batch normalization layer
@@ 4.3 Architecture-based
Reference (year) CL scenario & application CL technique
@@
Ozdemir and Goksel (2019) MRI based knee segmentation (2 Epi.) CIS Knowledge distillation + experience replay + multi-head
@@
Ravishankar et al. (2019) Chest x-ray view classification: DIS (2 Ep.), TIS (2 epi) Pneumothorax identification from X-ray: data-incremental (4 batches, each of 2K) Latent replay where samples are selected either randomly or importance based (farther from cluster centroids) + dynamic network (task specific feature transformer layers)
Lenga et al. (2020) Chest x-ray classification: DIS (2 epi) Comparative study: joint training, EWC, LwF
@@
Li et al. (2020) Dermoscopic images based skin disease classification: CIS (4-20 Ep.) Dual knowledge distillation (from old classifier (previous classes) and the fine-tuned classifier (a new FC layer fine-tuned on new classes)) + experience replay (fixed subset of data for each class is stored), during inference nearest-mean-of-exemplars method used for class prediction (Rebuffi et al. 2017) (where the mean is often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class)
Morgado et al. (2021) DIS: dermatological imaging modality classification Comparative study: Regularization (EWC), rehearsal based (AGEM, experience replay)
Derakhshani et al. (2022) TIS, CIS, DIS (4 epi) Regularization, memory-replay (comparative: EWC, MAS, LwF, iCaRL, EEIL)
Ranem et al. (2022) MRI based binary hippocampus segmentation: DIS (3 epi) Comparative study: transformers with EWC fisher matrix (Kirkpatrick et al. 2017), EWC with Riemannian walk (Chaudhry et al. 2018b), Modeling the Background (knowledge distillation and a modified Cross Entropy Loss) (Cermelli et al. 2020), Pseudo-labeling and Local Pod (multi-scale spatial distillation loss with pseudo labeling) (Douillard et al. 2021), Pooled Outputs Distillation (Douillard et al. 2020)
@@
Liu et al. (2022) CIS: CT data based segmentation of abdomen organs (4 Ep. to sequentially include organs) Rehearsal (latent replay) + regularization (knowledge distillation)
@@
Akundi and Sivaswamy (2022) Chest x-ray classification in CIS (5 epi) Regularization (knowledge distillation) + architecture (task specific head)
@@
Shu et al. (2022) DIS for fundus disease classification (2-4 epi) Replay (memory buffer optimization) + regularization (gradient regularization)
Kaustaban et al. (2022) DIS, CIS, TIS, Data-IL Regularization, replay (comparative study: EWC, LwF, CoPE, iCaRL, A-GEM)
@@
Chee et al. (2023) CIS in classification setting for three applications: cancer classification, diabetic retinopathy classification, skin lesions classification Architecture (separate high-level descriptor for each task) + regularization (distillation loss) + replay (data based rehearsal, fixed memory per task)
@@
Zhang et al. (2023b) (a) CIS for multi-organ segmentation, (b) CIS for abdomen segmentation to liver tumor segmentation Architecture: organ specific segmentation head + pseudo labelling based knowledge distillation
@@
Ji et al. (2023) CIS for multi-organ segmentation Architecture: organ specific decoder + pseudo labelling based knowledge distillation
@@
Zhang et al. (2023c) CIS: continual disease classification in different image modality settings Experience replay + Architecture based (keeps task specific adapters and classifier heads)
Bai et al. (2023) CIS Regularization: knowledge distillation
@@
Sadafi et al. (2023) DIS, CIS, DIS+CIS Experience replay based rehearsal (representativeness of examples for each class and model uncertainty) + regularization (binary-cross entropy based distillation loss)
@@
Yang et al. (2023b) MedMNIST based disease classification: CIS+DIS (4 epi) Generative replay+ data-focused regularization
@@
Bándi et al. (2023) DIS (cancer detection across organs such as breast, colon, and head-neck) Comparative study of existing approaches: PackNet (architecture), EWC (regularization), GEM (rehearsal)

### v4

@@ 4 Continual learning strategy
Owing to frequently encountered domain shifts in medical applications, CL strategies aim to tackle performance drop due to domain shifts and thus might become a necessity in clinical applications Pianykh et al. (2020); Lee and Lee (2020). All the available strategies for preventing catastrophic forgetting can be broadly categorized into three main categories: (a) rehearsal-based approaches where a small memory is used to store previous episode data in some form, (b) regularization-based methods where the aim is to control weight update to minimize forgetting the previous learning, (c) architectural-based methods which mainly aim to keep some network parameters isolated for each episode, and (d) hybrid category which offers various combination of any of these three categories. Figure 6 presents a taxonomy of CL techniques and their adaptation in the medical field by various works. Further, the sub-technique category for each work is also mentioned in the figure. More detail of CL techniques in each work can be found in  Table 10,  Table 11,  Table 12,  Table 13, and  Table 14 for rehearsal, regularization, architectural, hybrid, and comparative studies, respectively. Additionally, a pie chart showing the ratio of works falling in the above-mentioned categories is shown via Figure 5. It indicates the popularity of using one or more strategies together in hybrid settings owing to better classification or segmentation performances in sequential learning. Now, we discuss each category in detail in the following sections.
@@
Reference (year) Application & CL scenario CL technique
Hofmanninger et al. (2020a) Chest CT classification: DIS (3 Ep.) Experience replay with a dynamic memory size where samples are ranked by computing Gram matrix on the last activation maps
Srivastava et al. (2021) Chest X-ray classification: DIS(3 Ep.) Latent replay
Perkonigg et al. (2021) Brain age estimation: DIS (3 Ep.) Experience replay
Perkonigg et al. (2022) 3 separate experiments for 3 applications: (a) cardiac segmentation: DIS (4 Ep.), (b) lung nodule detection: DIS (4 Ep.), and (c) brain age estimation: DIS (4 Ep.) Experience replay
Karthik et al. (2022) MRI data for brain sclerosis lesions segmentation: DIS (8 Ep.) Random samples from each episode kept for experience replay
Li et al. (2022) CMR based Cardiac segmentation across 4 scanners: DIS (4 Ep.) Generative replay: style-oriented replay
Bera et al. (2023) (a) Binary segmentation of prostate, spleen, and hippocampus: TIS, (b) binary hippocampus segmentation across 2 datasets: DIS, (c) binary prostate segmentation across 4 datasets: DIS Experience replay based rehearsal strategy: store max K exemplars (half of the exemplars based on fraction of positive class content inside the image and other half based on gradient variation)
Wang et al. (2023b) (a) Segmentation for endoscopy: CIS (2-3 Ep.) (b) surgical instrument segmentation (2 Ep.) Generative replay
Byun et al. (2023) (a) Diabetic retinopathy severity classification: DIS (2 Ep.), (b) dermoscopy skin lesion detection: DIS (3 Ep.) Generative replay: The generator model is a conditional text-to-image diffusion model that is updated with current episode data and then used to generate labeled data for the current episode and stored in the buffer. Current episode data along with labeled data for each episode kept in the buffer, is utilized for fine-tuning the classifier.
Huang et al. (2023) Tumor subtype classification at WSI-level: CIS (4 Ep.) Experience replay
Sun et al. (2023b) Disease classification in time-series signals: DIS (10 time-steps) Experience replay (adaptive importance based replay)
Wei et al. (2023) Brain tumor segmentation: IIS (8 Ep.) Experience replay
Xiao et al. (2023) Skin disease classification with dermoscopic and clinical images: CIS (18 Ep.) Rehearsal (Memory buffer optimization: extension of GEM approach)
Li and Jha (2023) Disease classification on physiological signals: DIS (2 Ep.), CIS (2 Ep.), TIS (2-3 Ep.) (a) Experience replay: training loss based sample selection for the buffer, (b) generative replay: parametric (GMM) or non-parametric (KDE) models as a generator for past episodes.
Thandiackal et al. (2024) Histopathology colorectal tissue classification: DIS (3 Ep.) GAN-based generative replay
Bringas et al. (2024) Alzheimer’s disease stage identification using motion-sensor: IIS (2 Ep.,3 Ep.,4 Ep.) Rehearsal (A-GEM)
Zhu et al. (2024a) Histopathology WSI retrieval: CIS (4 Ep.) Rehearsal (experience replay: memory bank with reservoir sampling)
@@ 4.1.2 Latent replay-based
While experience replay requires storing the past samples in raw form, it can cause serious privacy violations in critical medical applications Thandiackal et al. (2024); Zhu et al. (2024b); Bayasi et al. (2024). Therefore, a less concerning direction, i.e., storing features instead of raw images, is also explored. For continual chest X-ray classification application in domain shift conditions, Srivastava et al. (2021) explore leveraging vector-quantization to store and replay hidden representations under memory constraints. Although not directly accessible, sharing of latent representations involves possible privacy threats through manipulation and reconstruction of actual sensitive medical data Pennisi et al. (2023). Given access to a model and its latent space, a malicious entity could reverse-engineer a patient’s medical record, violating privacy. Attackers might create adversarial instances to alter latent representations, potentially reconstructing crucial information from seemingly benign data. Reconstruction attacks Newaz et al. (2020), membership inference attacks Shokri et al. (2017), model inversion attacks Fredrikson et al. (2015), etc., can be potential challenges associated with sharing of features. Models like GAN, auto-encoder, variational auto-encoder, etc, inherently possess the capability to regenerate raw data from latent spaces. Appropriate privacy measures like differential privacy Abadi et al. (2016), privacy-preserved neural networks Jovanovic et al. (2022), adversarial training Yi et al. (2019), encryption and secure multi-party computation Spini et al. (2024), etc. need to be considered to protect latent features during processing and sharing.
@@ 4.1.4 Memory buffer optimization
Experience replay-based approaches might overfit the stored sub-samples and seem to be bounded by joint training. Alternatively, memory buffer optimization-based approaches following constrained optimization provide solutions that have more scope for both plasticity and stability. The core idea is to guide the training of the current episode by utilizing the stored buffer samples from the past. Typically, the gradients of the current task are projected such that they do not negatively impact the gradients computed on the buffer samples. Thus, it is ensured that the model performs well on all previous episodes while learning a new episode. For example, the Gradient Episodic Memory (GEM) Lopez-Paz and Ranzato (2017) approach corrects the gradient computed on a mini-batch during stochastic gradient descent by utilizing an exemplar sample set from past episodes. This prevents changes that could degrade the performance of the network on the exemplar set. Further, the A-GEM approach Chaudhry et al. (2018a) relaxes the projection in one direction computed by randomly selected samples from the memory buffer where both the buffer and sampling size are tunable hyper-parameters. There are some early attempts in this area in the medical realm. For example, for the fundus disease diagnosis problem, Shu et al. (2022) presents a gradient regularisation approach for preventing forgetting and a replay-oriented consistency calculation method combined with a subspace weighting strategy to promote model adaptability. The authors employ gradient regularisation in conjunction with a replay-oriented technique. The replay-oriented technique adaptively improves the update for incremental domains without carefully choosing exemplars for replay, whereas gradient regularisation preserves information from previous ones.
@@
Reference (year) Application & CL scenario CL technique
Baweja et al. (2018) MRI-based normal brain structures segmentation: TIS (2 Ep.) EWC
McClure et al. (2018) Axial and sagittal brain segmentation: DIS (4 Ep.) Distributed Weight Consolidation (DWC)
Zhang and Wang (2019b) Longitudinal MRI-based Alzheimer’s disease progression modeling: TIS (7 Ep.) EWC
van Garderen et al. (2019) MR data-based glioma segmentation: DIS (2 Ep.) EWC
Özgün et al. (2020) Brain MRI segmentation: DIS (4 Ep.) Learning rate regularization built over memory aware synapses technique
Chen and Tang (2022) Histopathology-based breast cancer classification: CIS (4 Ep.) Regularization: EWC
Zhang et al.; Zhang et al. (2023d; 2021) (a) MRI-based prostate segmentation: DIS, (b) retinal image-based optic cup and disc segmentation: DIS Shape and semantics-based selective regularization to penalize changes of parameters with high joint shape and semantics-based importance
Zhu et al. (2023) (a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites: DIS, (b) cross-site and cross-modality 4-class abdominal segmentation between CT and MRI scans: DIS, (c) same-site cross-modality muscle segmentation of muscles and bones between MRI and mDixon scans: DIS Knowledge distillation
Liu et al. (2023a) Brain tumor segmentation: CIS with domain shift (3 Ep.) Pseudo label-based knowledge distillation
Bai et al. (2023) Surgical question answering: CIS Regularization: knowledge distillation
Wu et al. (2024) Image super-resolution: TIS (4 Ep.) Regularization (knowledge distillation and parameter importance based gradient update)
Aslam et al. (2024) Disease outbreak detection in time-series signal: DIS (10 Ep.) Regularization (EWC)
Zhu et al. (2024b) (a) MRI Prostate segmentation: DIS (b) MRI Cardiac segmentation: DIS Regularization (knowledge distillation)
@@ 4.2.1 Prior-focused regularization
Apart from EWC and SI, various other methods have been contributed. Popularly, the Memory Aware Synapses (MAS) approach Aljundi et al. (2017) calculates the importance of weights with a model of Hebbian learning in the biological system, which relies on the sensitivity of the output function and can hence be utilized in an unsupervised manner. For medical segmentation, a MAS-inspired learning rate regularization approach was contributed by Özgün et al. (2020) for sequential training on different domains. It involves reducing the learning rate for important parameters to prevent forgetting, offering a more direct approach compared to the surrogate loss used in MAS. Yet another novel emerging direction is Orthogonal Weight Modification (OWM) Zeng et al. (2019), where the gradient of the current task is projected into the orthogonal direction to the subspace spanned by gradients of all previous tasks. Then parameter updates for the current task are permitted only in the orthogonal direction of the past episodes, thus protecting against interference with already learned knowledge.
@@ 4.2.2 Data-focused regularization
The data-focused regularization method Li and Hoiem (2017) aims to distill knowledge from a model trained on the previous tasks to the model trained on the new task in order to consolidate previously learned knowledge. Typically, the previous model acts as a teacher and the current model as a student while adopting knowledge-distillation (KD) Gou et al. (2021) to avoid catastrophic forgetting. Usually, all the old data should be available in knowledge-distillation, which is not the case with CL; therefore, as an alternative, few old data, current data, or the generated old data are explored for the same. The loss function has an additional distillation loss for replayed data. Each input is replayed with a soft target obtained using the stored model.
Silver and Mercer (2002) first proposed to use previous task model outputs given new task input images, mainly for improving new task performance. Later, in natural image classification application, Li and Hoiem (2017) re-introduced the concept as the LwF technique. A copy of the previous model parameters is stored before learning the new task, and then it is used to get the soft labels for the new task as the target from the classifiers of the previous tasks. The available ground truth is used as the target for the new task classifier. LwF has been adopted in various medical applications like incremental brain MRI segmentation Ozdemir et al. (2018), chest X-ray classification Lenga et al. (2020), disease classification Derakhshani et al. (2022), histopathology tumor classification Kaustaban et al. (2022), etc.
Other works in the natural image processing domain Jung et al. (2016); Zhang et al. (2020) have been introduced with LwF-related ideas; however, it has been shown that this strategy is vulnerable to domain shift between tasks Aljundi et al. (2017). It was pointed out that LwF would not result in good performance if the data distributions between different tasks are quite diverse Rannen et al. (2017). To overcome this, Rannen et al. (2017) facilitate incremental integration of shallow auto-encoders to constrain task features in their corresponding learned low dimensional space. Hence, they further trained an autoencoder for each task to learn the most important features corresponding to the task and used it to preserve knowledge. Such multiple expert-based knowledge-distillation have proven effective in domain generalization application too Niu et al. (2023). Student expert modules possess domain-specific information. Each expert also learns from all other students by knowledge-distillation technique to facilitate domain-invariant features. Also, when more emphasis is attributed toward improving unseen classes, i.e., forward transfer, CL based sequential learning of domains helps in domain generalization Li et al. (2020b).
@@
Reference (year) Application & CL scenario CL technique
@@
Bayasi et al. (2021) Skin lesion image classification: DIS (6 Ep.) Architecture (fixed but partitioned network: culprit unit pruning mechanism)
@@
González et al. (2022) MRI-based Hippocampus Segmentation: DIS (2 Ep.) Architecture based approach based on out-of-distribution detection concept: maintain multivariate Gaussians for all past learning created on batch normalization layer
Yang et al. (2023b) Skin disease classification on dermoscopic, clinical images: CIS (4-20 Ep.) Pretrained feature extractor + class specific GMMs built on each deep feature
Verma et al. (2024) Disease classification on (a) Fundus: TIS (2 Ep.), (b) pathology images: TIS (3 Ep.) Architectural (Task specific model is trained, task id is inferred during inference)
Qazi et al. (2024) (a) Disease classification in histopathology images: CIS (7 Ep.), (b) Skin lesion classification: CIS (4 Ep.) Task specific adapter with merging facility to increase computational efficiency
@@ 4.3 Architecture-based
Reference (year) Application & CL scenario CL technique
@@
Ozdemir and Goksel (2019) MRI based knee segmentation: CIS (2 Ep.) Knowledge distillation + experience replay + multi-head
@@
Ravishankar et al. (2019) Chest X-ray view classification: DIS (2 Ep.), TIS (2 Ep.), Pneumothorax identification from X-ray: IIS (4 batches, each of 2K) Latent replay where samples are selected either randomly or importance based (farther from cluster centroids) + dynamic network (task specific feature transformer layers)
@@
Li et al. (2020a) Dermoscopic images based skin disease classification: CIS (4-20 Ep.) Dual knowledge distillation (from old classifier (previous classes) and the fine-tuned classifier (a new FC layer fine-tuned on new classes)) + experience replay (fixed subset of data for each class is stored), during inference nearest-mean-of-exemplars method used for class prediction (Rebuffi et al. 2017) (where the mean is often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class)
@@
Liu et al. (2022) CT data based segmentation of abdomen organs: CIS (4 Ep.) Rehearsal (latent replay) + regularization (knowledge distillation)
@@
Akundi and Sivaswamy (2022) Chest X-ray classification: CIS (5 Ep.) Regularization (knowledge distillation) + architecture (task specific head)
@@
Shu et al. (2022) Fundus disease classification: DIS (2-4 Ep.) Replay (memory buffer optimization) + regularization (gradient regularization)
@@
Chee et al. (2023) 3 applications as (a) Cancer classification: CIS, (b) diabetic retinopathy classification: CIS, (c) skin lesions classification: CIS Architecture (separate high-level descriptor for each task) + regularization (distillation loss) + replay (data based rehearsal, fixed memory per task)
@@
Zhang et al. (2023b) (a) Multi-organ segmentation: CIS, (b) abdomen segmentation to liver tumor segmentation: CIS Architecture: organ specific segmentation head + pseudo labelling based knowledge distillation
@@
Ji et al. (2023) Multi-organ segmentation: CIS Architecture: organ specific decoder + pseudo labelling based knowledge distillation
@@
Zhang et al. (2023c) Continual disease classification in different image modality settings: CIS Experience replay + Architecture based (keeps task specific adapters and classifier heads)
@@
Sadafi et al. (2023) Dermoscopic WBC classification: DIS, CIS, DIS+CIS Experience replay based rehearsal (representativeness of examples for each class and model uncertainty) + regularization (binary-cross entropy based distillation loss)
@@
Yang et al. (2023c) MedMNIST based disease classification: CIS+DIS (4 Ep.) Generative replay+ data-focused regularization
Hua et al. (2023) sEMG-based Gesture Classification: CIS (4 Ep.) Rehearsal (experience replay) and Regularization (KD)
Chen et al. (2023) 3 separate applications including (a) Optic disc: DIS (3 Ep.), (b) cardiac: DIS (3 Ep.), (c) prostate segmentation: DIS (6 Ep.) Rehearsal (GAN based generative replay) + Regularization (KD from past segmentation model)
Sun et al. (2023a) Multi-class classification on time-series signals: CIS (4-10 Ep.) Rehearsal (experience replay)+ Regularization (KD)
Li et al. (2024) MR Cardiac segmentation: DIS (5 Ep.) Rehearsal (privacy-aware generative replay with CGAN) + Architecture (ConvLSTM based domain-customised expansion block)
Ceccon et al. (2024b) Chest X-ray disease classification: NIC (7 Ep.) Pseudo-Labeling and memory buffer based replay
Bayasi et al. (2024) (a) Skin lesion classification: CIS (3 Ep.), DIS (4 Ep.), CIS+DIS (5 Ep.), (b) blood cell classification: CIS (4 Ep.), (c) Colon tissue classification: CIS (4 Ep.) Architecture (fixed but partitioned network: culprit unit pruning mechanism) + Regularization (knowledge distillation from a generalized network to expert network to enhance generalizability)
Ye et al. (2024) Multi-modal (medical report, MRI, X-ray, CT, histopathology) Representation learning: TIS (5 Ep.) Rehearsal (K-means sampling based buffer creation for experience replay ) + Regularization (MSE based knowledge distillation)
Table 14: List of various comparative studies of CL techniques in literature
Reference (year) Application & CL scenario CL technique
Lenga et al. (2020) Chest X-ray classification: DIS (2 Ep.) Comparative study: joint training, EWC, LwF
Morgado et al. (2021) Dermatological imaging modality classification: DIS Comparative study: Regularization (EWC), rehearsal based (AGEM, experience replay)
Derakhshani et al. (2022) Disease classification: TIS, CIS, DIS (4 Ep.) Regularization, memory-replay (comparative: EWC, MAS, LwF, iCaRL, EEIL)
Ranem et al. (2022) MRI-based binary hippocampus segmentation: DIS (3 Ep.) Comparative study: transformers with EWC fisher matrix (Kirkpatrick et al. 2017), EWC with Riemannian walk (Chaudhry et al. 2018b), Modeling the Background (knowledge distillation and a modified Cross Entropy Loss) (Cermelli et al. 2020), Pseudo-labeling and Local Pod (multi-scale spatial distillation loss with pseudo labeling) (Douillard et al. 2021), Pooled Outputs Distillation (Douillard et al. 2020)
Kaustaban et al. (2022) Tumor classification: DIS, CIS, TIS, IIL Regularization, replay (comparative study: EWC, LwF, CoPE, iCaRL, A-GEM)
@@
Bándi et al. (2023) Cancer detection across organs such as breast, colon, and head-neck: DIS Comparative study of existing approaches: PackNet (architecture), EWC (regularization), GEM (rehearsal)
Verma et al. (2023) Disease classification on (a) Fundus: CIS (3 Ep.), (b) pathology images: CIS (3 Ep.) Comparative study of existing buffer-free (privacy-aware) approaches in three categories: Rehearsal (GR, BIR), Regularization (EWC, SI, MAS, MUC-MAS, RWalk, OWM, GPM, LwF, LwM), Architecture (EFT)
Kim et al. (2024) Arrhythmia detection on ECG data: DIS (4 Ep.) Comparative study of 3 existing regularization techniques (LwF, EWC, MAS)
Ceccon et al. (2024a) Chest X-ray disease classification: CIS (5 Ep.) Comparative study of approaches in regularization (LwF, pseudo-label), rehearsal (replay), hybrid (rehearsal+regularization) categories for fairness evaluation.